A company's stock recently paid a dividend of $4.02 and expects dividends to grow annually at a rate of 6.2%. Based on the CAPM, the required return for this stock is estimated to be 13.0%. Based on this, the maximum that you would be willing to pay for the stock is $__.__. Round your answer to 2 decimal places.
Answer #1

Ans $ 62.78

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
P0 = 4.26924 / (13%- 6.2%)
P0 = 62.78
D1 = D0* (1 + g)
D1 = 4.02* (1 + 6.2%)
D1 = 4.26924